OpenStudy (anonymous):

Solve: 3(x + 4) = 5(x − 2) 11 −11 1 −1

OpenStudy (anonymous):

3(x + 4) = 5(x − 2) multiply to start 3x+12=5x-10 now combine like terms by adding 10 to both sides 3x+22=5x now you must subtract 3x from both sides to get X alone 22=2x now divide both sides by 2 X=11

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

no problem